OpenStudy (anonymous):

Did I do this Algebra problem correctly? 8(8.8 + 10^12) Simplify the expression. Write using the scientific expression. Is it 7.04 x 10^13

OpenStudy (anonymous):

no 8(8.8 + 10^12) = 8*8.8 + 8*10^12

OpenStudy (anonymous):

What?Oh it's 10^12

OpenStudy (anonymous):

Sorry! Am I right? :D

OpenStudy (anonymous):

70.4 + 8 * 10^13

OpenStudy (anonymous):

ops ^12 not 13 ._.\[8(8.8 + 10^{12})\]\[70.4 + 8 \times 10^{12}\]